A Complete Bibliography of Philosophy of Science: 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 September 2017 Version 1.03

Title word cross-reference

1934 [117]. 1937 [181]. 1938 [276].

Abel [311]. Abraham [79]. Absolute [299]. Absolute-Relative [299]. Abstract [44, 176]. Abstraction [287]. Action [172, 309]. Age [26]. Aim [251]. Aims [75]. Alban [304]. Albert [307]. Alexander [72]. Alfred [143, 226]. Algebra [158]. Ambiguity [271]. America [114]. among [26]. Analysis [196, 57, 218, 47]. Anatomy [189]. Ancient [26]. Anderson [227]. Announcement [82, 266, 282, 313]. Announcements [95]. Anthropological [290]. Anwendung [191]. Application [9]. Applications [170]. Approach [251]. Apriori [131]. Argument [134]. Arguments [246]. Aristotelian [64, 17]. Aristotle [156, 79]. Arne [212, 301]. Arranging [174, 244]. Arrow [227]. Arthur [26]. Aspects [103, 46]. Assertion [39, 23]. Atom [284, 293]. Attributes [74, 108]. Auflage [191]. Avenir [154].

B [308]. Babel [237]. Back [14, 29, 42, 55]. Basic [101]. Basis [7]. Bavink [11]. Baylis [307]. Bearing [220]. Behanan [201]. Being [279]. Belief [192]. Belin [326]. Bell [66]. Beneze [280]. Benjamin [92]. Bennett [307]. Bernhard [11]. between [121]. Beyond [32, 325]. Biological [220]. Biology [89, 311]. Biophysics [20, 61]. Biser [298]. Black [65]. Boll [118]. Book [83, 65, 11, 240, 278, 302, 327, 212, 79, 275, 279, 191, 305, 179, 165, 143, 54, 92, 96, 182, 166, 276, 277, 299, 301, 303, 304, 306, 307, 308, 309, 310, 311, 324, 325, 326, 80, 91, 118, 144, 141, 189, 227, 298, 280, 281, 117, 154, 225, 226, 264, 66, 156, 155, 192, 201, 213, 300, 26, 40, 239, 25, 53, 93, 10, 142, 190, 78]. Books [94, 81, 119, 132, 145, 157, 180, 214, 229, 241, 265, 312, 276]. Border [271]. Border-Line [271]. Boyer [308]. Bridgman [141]. Britton [327]. Buros [276].

C [305, 165, 92, 78]. C. [165]. Calculus [308]. Caricature [118]. Carl [308]. Case [83, 54, 96, 166]. Case-Book [83, 54, 96, 166]. Cases [271]. Cassirer [225]. Causal [87, 317]. Causality [184, 16, 220]. Causation [220]. Century [162]. Cercle [154]. Chance [155, 155]. Chapter [171]. Character [2, 3]. Charles [165, 307, 300]. Chemistry [37, 311]. Chimiques [281]. Clarence [25]. Clarification [318]. Classification [27, 41, 178]. Cohen [53, 227]. Coincidences [31]. Collected [165, 165]. Comment [164]. Common [234]. Communication [327]. Communist [58]. Complementarity [184]. Complementary [163]. Conceived [301]. Concept [203, 89, 222, 129, 72]. Conception [211, 161, 26]. Concepts [308, 101, 142]. Conceptual [210]. Concerning [63, 323, 297, 17, 272]. Congress [228]. Connected [170]. Considered [298]. Consistency [126]. Contemporary [263]. Continued [160, 41, 153]. Continuity [151, 150]. Continuum [192]. Contradiction [272]. Contribution [47]. Contributions [203]. Cooper [25]. Coordinate [31]. Corrections [181]. Cosmology [78]. Criterion [122]. Critical [203, 186, 318, 65]. Criticism [263]. Critique [89, 197, 280]. Crystalline [310]. Cultural [292]. Culture [223, 291, 261].

D [304]. Dangers [170]. Danish [251]. Debate [58]. Deduction [123]. Definitons [281]. Degrees [231]. Denker [275]. DesCartes [194]. Design [136, 107]. Detail [127]. Determinismus [225]. Development [48]. Developpement [154]. Dewey [324, 264]. Dialectics [58]. Dialektik [10]. Dimensional [300]. Dimensions [74]. Directive [172]. Discovery [22]. Doctrine [140, 153]. Doubt [194]. Drei [278]. Dualism [274]. Dubs [295]. D’Une [326]. Durch [302]. D’Utilite [326].

Economic [140, 153]. Edel [79]. Editorial [217, 1]. Einfuhrung [191, 226]. Einheit [117, 117]. Einstein [182]. Embedding [198]. Emergence [255, 256]. Emile [208]. Emotion [322]. Empirical [135, 164, 294, 122, 191]. Empirisme [154]. Encyclopaedia [175]. Encyclopaedism [251]. Engels [10]. England [162]. Enstehung [302]. Enzymes [310]. Eric [66].

Erkenntnis [212]. Ernest [53]. Ernst [225]. Errata [215]. Especial [311]. Essai [281]. Essays [143]. Evaluation [201]. Evolution [269, 114]. Examination [179, 6, 19, 45, 88, 246]. Excluded [272]. Exercise [196]. Experience [108, 90, 240]. Experiment [174, 244]. Experimentales [281]. Explanation [208]. Extension [231]. External [294, 238, 134, 221]. Extra [296]. Extra-Sensory [296]. Ezra [227].

Fact [209, 320, 125]. Facts [235]. Fairies [219]. Faiths [303]. False [124]. Fare [144]. Feeling [40]. Field [220]. Figures [275]. Finite [4]. First [171]. Flexibility [76, 52, 38]. Force [326]. Forgotten [75]. Formal [307]. Forms [108]. Foundations [181, 167, 20, 5, 144]. Fourth [228]. Francaise [326]. Franklin [299]. Freedom [231]. Freud [35]. Friedrich [10]. Front [15, 30, 43, 56, 67, 84, 97, 106, 120, 133, 147, 159, 168, 183, 193, 202, 216, 230, 242, 254, 267, 283, 314, 329]. Fruitfulness [123]. Fulton [91]. Function [73, 298]. Future [249].

G [182, 80, 280]. Galileo [156]. Geist [279]. Gene [315]. General [298]. Genetics [34]. George [309]. Gestalt [46, 21]. Gestalt-Problem [46]. Gestalten [275]. Given [69]. God [239]. Greeks [26]. Growth [73]. Gruenberg [92].

H [11, 182, 309, 310]. Haeckel [34]. Hailperin [181]. Haldane [50]. Harold [25]. Hartshorne [165]. Hatfield [11]. Heidel [26]. Herman [279]. Heroic [26]. Histoire [190]. History [277, 261, 110, 190, 277]. Hoenigswald [275]. Home [24]. Hooke [252]. Hormones [311]. Hudson [303]. Human [100]. Humanizing [142]. Hyman [227]. Hypothesis [274, 104].

Idealism [296]. Ideals [26]. Idee [326]. Idee-Force [326]. If [258]. Ignaz [302]. Ihre [191]. II [112, 19, 244]. III [146, 45]. Immanuel [306]. Immediate [90]. Immunity [131]. Imperatives [321]. Implications [210, 322]. Impossibility [121]. Imputation [317]. Independence [126]. Indeterminism [8]. Indeterminisms [207]. Indeterminismus [225]. Indian [210]. Indices [292]. Indifference [295]. Induction [51, 294, 3, 238, 116, 204, 221]. Inductive [134]. Inexact [130]. Infinite [4, 211, 79]. Information [105, 253, 328]. Inquiry [264]. Instances [152]. Insulin [311]. Interaction [121]. International [228]. Internationalen [117]. Interpretation [224, 285]. Introduction [277, 53]. Investigation [317]. Irving [25]. Irwin [298]. Isaac [239]. Italian [275]. Italienischen [275]. IV [88].

J [278, 277, 299, 325, 91]. Jacob [311]. James [277]. January [181]. Jay [303]. Jean [326]. Johannes [278]. John [310, 311, 324]. Julius [179].

Kant [306]. Karl [327]. Knowledge [62, 288, 273, 49, 316]. Kongresse [117]. Kovoor [201]. Kronecker [127]. Kurt [156].

L [182, 142]. Landford [25]. Language [210, 198]. Languages [292]. Law [102, 222, 115]. Laws [291, 272]. Leaning [156]. Lebens [302]. Lee [325]. Leibniz [128, 50]. Less [176]. Lewin [156]. Lewis [25]. Lichtig [302]. Lieber [182]. Life [172, 285, 213]. Light [34]. Line [271]. Linguistic [103]. Location [77]. Lodestone [162]. Logic [181, 263, 259, 167, 270, 64, 321, 17, 9, 185, 205, 137, 198, 44, 53, 188, 24, 110, 307, 25, 264]. Logical [149, 196, 47, 3, 179, 189, 264]. Logics [260, 176]. Logik [226]. Logique [326, 118, 154]. Logistic [93]. Lost [139].

M [300]. Mannoury [80]. Many [9, 260, 231]. Many-Valued [9, 260]. Marcel [118]. Mass [262, 297]. Matchette [299]. Materialistic [285]. Mathematical [181, 161, 167, 20, 46, 100, 170]. Mathematics [65, 248]. Mathematische [226]. Matter [14, 29, 42, 55, 15, 30, 43, 56, 67, 84, 97, 106, 120, 133, 147, 159, 168, 183, 193, 202, 216, 230, 242, 254, 267, 283, 314, 329, 162, 121, 128, 99]. Max [65]. Meaning [135, 177, 164, 39, 148, 160, 23, 122, 169, 16]. Meanings [223]. Measurement [171]. Mechanical [161]. Mechanics [58, 7, 269, 129, 50, 305]. Mechanism [104, 32]. Medicine [311]. Memory [192]. Mesure [280]. Metaphysics [232, 86, 165, 107, 165]. Method [194, 31, 35, 18, 58, 63, 48, 199, 36, 231, 53]. Methodology [276, 60, 71]. Methods [70, 112, 146, 26]. Meyerson [208]. Middle [272]. Miller [224]. Mind [121, 279, 92, 80]. Minds [288]. Mises [191]. Modal [188]. Modality [44]. Modern [60, 71, 186, 250, 142]. Modernen [225]. Modes [298]. Morris [53]. Mr. [224]. Mystery [22, 256].

Nagel [53]. Natur [10]. Natural [11, 289, 8, 298, 292]. Nature [65, 57, 161, 248, 108, 245, 141, 298, 69, 234, 286]. Neither [124]. Ness [212, 301]. Neubearbeitete [191]. Neue [191]. Neurath [154]. Newton [239]. nor [124]. North [143]. Northrop [310]. Note [123, 64, 111]. Notes [83, 54, 96, 166]. Notices [13, 28]. Number [63]. Number-Pairs [63].

O [276, 325]. Objective [215, 173]. Objectivity [187]. Observer [138]. Ockham [151]. Old [303, 239]. Oliver [142]. Operational [218]. Operationalism [197]. Operations [141, 281]. Order [236]. Organismic [104, 49, 110]. Organization [62, 33]. Orman [93]. Other [288, 311]. Otto [154]. Our [288]. Outline [100]. Outlines [135, 164].

P [141]. Pairs [63]. Papers [165]. Paradoxer [12]. Paradoxers [27, 41]. Parallel [48]. Parapschologie [278]. Parsimony [111]. Pattern [163]. Patterns [261]. Paul [165, 281]. Peace [306]. Pedagogical [251]. Peirce [165, 165]. Pensee [190]. Perception [296]. Perish [303]. Perpetual [306]. Perry [300]. Philosophers [301]. Philosophic [57, 2]. Philosophical [322, 129, 190, 143]. Philosophie [278]. Philosophy [247, 263, 68, 7, 217, 1, 85, 205, 258, 40, 233, 91, 142]. Physical [77, 31, 150, 99, 141, 186, 136, 258]. Physicists [219]. Physico [46]. Physico-Mathematical [46]. Physics [259, 18, 287, 48, 197, 169, 249, 144, 9, 60, 71, 260, 220, 171, 239, 144]. Physik [225]. Place [290]. Points [186]. Poles [80]. Poortman [278]. Positive [149]. Positivism [47, 179]. Positivists [86]. Possible [68]. Postulates [199]. Postulational [126, 70, 112, 146]. Potentiality [109]. Prager [117]. Pragmaticism [165]. Pragmatism [165]. Preamble [49]. Preconceptions [101]. Prediction [240]. Preface [245]. Preview [299]. Principle [51, 3, 38, 295]. Principles [305, 156]. Probabilities [5]. Probability [181, 240, 203, 259, 294, 167, 191, 260, 273, 129, 116, 204]. Problem [77, 47, 99, 33, 46, 75]. Probleme [275]. Problems [2, 275]. Procedure [291]. Process [264]. Production [262, 297]. Professional [301]. Professor [295]. Program [127]. Progress [211]. Propaedeutic [113]. Proposal [39, 23]. Proposition [125]. Propositions [124, 158]. Pseudo [140, 153]. Pseudo-Scientific [140, 153]. Psychical [250]. Psycho [77]. Psycho-Physical [77]. Psychology [215, 173, 48, 192, 250, 163, 40, 113, 90, 156]. Psychophysical [47]. Public [92]. Purpose [136, 107]. Purposes [317].

Quantum [7, 6, 19, 45, 88, 129, 50]. Quest [139]. Quine [93].

R [182, 53]. Rational [286]. Re [246]. Re-Examination [246]. Real [155]. Realism [318, 246, 300]. Realistic [263]. Reality [293]. Received [94, 81, 119, 132, 145, 157, 180, 214, 229, 241, 265, 312]. Reference [311]. Reiser [142]. Relation [74, 311, 101]. Relations [215, 173, 100]. Relative [299]. Relativity [136, 182, 78]. Renaissance [275]. Renaud [281]. Renegade [152]. Research [262, 297, 291, 250, 276]. Review [81, 65, 11, 240, 278, 302, 327, 212, 79, 275, 279, 191, 305, 179, 165, 143, 92, 182, 276, 277, 299, 301, 303, 304, 306, 307, 308, 309, 310, 311, 324, 325, 326, 80, 91, 118, 144, 141, 189, 298, 280, 281, 117, 154, 225, 226, 264, 66, 156, 155, 192, 201, 213, 300, 26, 40, 239, 25, 53, 93, 10, 142, 190, 78]. Reviews [276, 227]. Revision [188]. Revolution [232, 326]. Richard [275, 191, 305, 78]. Rise [114]. Robert [252]. Role [138]. Rudolph [179].

S [165]. sa [118]. Samuel [72]. Sanders [165]. Scale [31]. Schema [298]. Schmalenbach [279]. Schools [163]. Schopfung [302]. Science [200, 215, 228, 247, 319, 103, 232, 173, 102, 289, 237, 217, 1, 189, 36, 101, 236, 85, 175, 137, 258, 156, 26, 44, 75, 113, 233, 234, 142, 110, 243, 91, 142, 92, 189]. Sciences [187, 62, 222, 174, 244, 11]. Sciencing [243]. Scientific [22, 35, 68, 58, 76, 208, 165, 52, 16, 36, 140, 153, 231, 316, 114, 201, 53, 38]. Scientifique [190]. Scientist [309, 268]. Search [66]. Sein [279]. Selection [8, 298]. Selectivity [298]. Sensation [74, 40]. Sense [234]. Sensory [296].

Sentence [209]. Seventeenth [162]. Sheen [91]. Shotwell [277]. Significance [129]. Silverberg [304]. Simples [98]. Simplicity [169]. Social [187, 102, 289, 222, 36, 101, 236]. Sociale [326]. Society [227]. Socrates [304]. Solid [144]. Some [124, 205, 322, 170, 246]. Sovereignty [227]. Space [108, 269, 61, 21, 72]. Space-Time [269, 72]. Span [213]. Spatial [77]. Species [89]. Spirit [114]. Stability [87]. Stafford [11]. Statistical [305, 276]. Statistics [178]. Statistik [191]. Status [16]. Stetige [302]. Stimulus [74]. Structure [257, 73, 323, 189, 209, 125]. Studies [290]. Study [101]. Supplement [181]. Survey [203, 65]. Symbolic [198, 25]. Symbolism [233]. Symposium [200]. System [62, 93]. Systems [87, 298, 9, 231].

T [66, 201]. Tarski [226]. Technique [126]. Techniques [101]. Technology [137]. Telepathy [250]. Testability [148, 160]. Their [101]. Theoretical [18, 249]. Theories [6, 19, 45, 88, 227, 269, 322]. Theory [255, 135, 164, 151, 79, 128, 290, 99, 245, 182, 186, 273, 264, 49, 100, 206, 5, 299, 141, 324]. There [98]. Thermodynamics [78]. Theses [205]. Things [109]. Thinkers [275]. Those [301]. Tide [213]. Time [37, 108, 269, 61, 21, 213, 171, 72, 206, 227]. Tolman [305, 78]. Tom [304]. Topological [156]. Topology [59]. Tower [237]. Towers [156]. Transitivity [115]. Tribute [143]. Troubled [65]. Troublesome [195]. True [124]. Truth [76, 301, 52, 38, 66]. Two [80]. Types [298].

Ultimate [98]. Understanding [162]. Unified [200, 99, 175]. Unity [228]. Universe [139, 141]. upon [220].

V [165]. Vagueness [319, 196, 271, 270]. Validity [64, 17]. Valuation [324]. Value [130, 320]. Valued [9, 260]. Values [235]. Verhalten [212]. Verifiability [235]. Verification [68]. versus [58, 161, 90]. VI [165]. Vicious [287]. Vienne [154]. Views [34]. Virtue [268]. Vital [33]. Vitalism [104, 32]. Vivisection [264]. Vol [165]. Volume [105, 253, 328]. Vorkonferenz [117]. Vortrage [278].

W [141]. Wahrheit [191]. Wahrscheinlichkeit [191]. Wahrscheinlichkeitslehre [191]. Wait [213]. Was [304]. Waters [65]. Weinberg [179]. Weiss [165, 165]. Where [217]. Whitehead [143, 224, 143]. Who [301, 304]. Wholes [257, 323]. Will [213]. Willard [93]. William [303, 309, 26, 151]. Winspear [304]. Wissenschaft [117]. Wissenschaftliches [212]. without [256, 107]. Woods [227]. World [240, 294, 238, 134, 155, 221].

X [195].

Yoga [201, 201]. Yonder [325].

Zweite [191].

References

Malisoff:1934:EWP

[1] W. M. Malisoff. Editorial: What is philosophy of science? Philosophy of Science, 1(1):1–4, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184478.

Carnap:1934:CPP

[2] Rudolf Carnap. On the character of philosophic problems. Philosophy of Science, 1(1):5–19, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184479.

Feigl:1934:LCP

[3] Herbert Feigl. The logical character of the principle of induction. Philosophy of Science, 1(1):20–29, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184480.

Bell:1934:FI

[4] E. T. Bell. Finite or infinite? Philosophy of Science, 1(1):30–49, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184481.

Struik:1934:FTP

[5] D. J. Struik. On the foundations of the theory of probabilities. Philosophy of Science, 1(1):50–70, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184482.

Malisoff:1934:EQTa

[6] William Marias Malisoff. An examination of the quantum theories. I. Philosophy of Science, 1(1):71–77, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184483.

Haldane:1934:QMB

[7] J. B. S. Haldane. Quantum mechanics as a basis for philosophy. Philosophy of Science, 1(1):78–98, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184484.

Fisher:1934:INS

[8] R. A. Fisher. Indeterminism and natural selection. Philosophy of Science, 1(1):99–117, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184485.

Margenau:1934:AMV

[9] H. Margenau. On the application of many-valued systems of logic to physics. Philosophy of Science, 1(1):118–121, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184486.

Struik:1934:BRB

[10] Dirk Jan Struik. Book review: Dialektik der Natur by Friedrich Engels. Philosophy of Science, 1(1):122–123, January 1934. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184487.

Benjamin:1934:BRB

[11] A. Cornelius Benjamin. Book review: The Natural Sciences by Bernhard Bavink; H. Stafford Hatfield. Philosophy of Science, 1(1):123–129, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184488.

M:1934:NP

[12] W. M. M. The new paradoxer. Philosophy of Science, 1(1):130, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184489.

Anonymous:1934:Na

[13] Anonymous. Notices. Philosophy of Science, 1(1):131, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184490.

Anonymous:1934:BMa

[14] Anonymous. Back matter. Philosophy of Science, 1(1):??, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184491.

Anonymous:1934:FMa

[15] Anonymous. Front matter. Philosophy of Science, 1(1):??, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184477.

Margenau:1934:MSS

[16] Henry Margenau. Meaning and scientific status of causality. Philosophy of Science, 1(2):133–148, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184385.

Kattsoff:1934:CVA

[17] L. Kattsoff. Concerning the validity of Aristotelian logic. Philosophy of Science, 1(2):149–162, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 184386. See note [64].

Einstein:1934:MTP

[18] Albert Einstein. On the method of theoretical physics. Philosophy of Science, 1(2):163–169, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184387.

Malisoff:1934:EQTb

[19] William Marias Malisoff. An examination of the quantum theories. II. Philosophy of Science, 1(2):170–175, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184388.

Rashevsky:1934:FMB

[20] N. Rashevsky. Foundations of mathematical biophysics. Philosophy of Science, 1(2):176–196, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184389.

Reiser:1934:TSG

[21] Oliver L. Reiser. Time, space and gestalt. Philosophy of Science, 1(2):197– 223, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184390.

Benjamin:1934:MSD

[22] A. Cornelius Benjamin. The mystery of scientific discovery. Philosophy of Science, 1(2):224–236, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184391.

Dewey:1934:MAP

[23] John Dewey. Meaning, assertion and proposal. Philosophy of Science, 1(2):237–238, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184392.

Weiss:1934:HL

[24] Paul Weiss. A home for logic. Philosophy of Science, 1(2):238, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184393.

Smith:1934:BRBa

[25] Henry Bradford Smith. Book review: Symbolic Logic by Clarence Irving Lewis; Cooper Harold Landford. Philosophy of Science, 1(2):239–246, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184394.

Robinson:1934:BRB

[26] Edward Schouten Robinson. Book review: The Heroic Age of Science. The Conception, Ideals, and Methods of Science among the Ancient Greeks by William Arthur Heidel. Philosophy of Science, 1(2):246–248, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184395.

M:1934:CP

[27] W. M. M. Classification of paradoxers. Philosophy of Science, 1(2):249– 250, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184396.

Anonymous:1934:Nb

[28] Anonymous. Notices. Philosophy of Science, 1(2):251, April 1934. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184397.

Anonymous:1934:BMb

[29] Anonymous. Back matter. Philosophy of Science, 1(2):??, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184398.

Anonymous:1934:FMb

[30] Anonymous. Front matter. Philosophy of Science, 1(2):??, April 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184384.

Bender:1934:MPC

[31] Wm. Bender. The method of physical coincidences and the scale coordinate. Philosophy of Science, 1(3):253–272, July 1934. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184589.

Singer:1934:BMV

[32] Edgar A. Singer, Jr. Beyond mechanism and vitalism. Philosophy of Science, 1(3):273–295, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184590.

Lillie:1934:PVO

[33] Ralph S. Lillie. The problem of vital organization. Philosophy of Science, 1(3):296–312, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184591.

Muller:1934:VHL

[34] H. J. Muller. The views of Haeckel in the light of genetics. Philosophy of Science, 1(3):313–322, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184592.

Brown:1934:FSM

[35] J. F. Brown. Freud and the scientific method. Philosophy of Science, 1 (3):323–337, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184593.

Mayer:1934:SMS

[36] Joseph Mayer. Scientific method and social science. Philosophy of Science, 1(3):338–350, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184594.

Hoagland:1934:CT

[37] Hudson Hoagland and Oliver L. Reiser. The chemistry of time. Philosophy of Science, 1(3):351–353, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184595.

Zwicky:1934:PFS

[38] F. Zwicky. On the principle of the flexibility of scientific truth. Philosophy of Science, 1(3):353–358, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184596.

Carnap:1934:MAP

[39] R. Carnap. Meaning, assertion and proposal. Philosophy of Science, 1(3): 359–360, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184597.

Ruddick:1934:BRF

[40] Chester Townsend Ruddick. Book review: The feeling of feeling: The Philosophy and Psychology of Sensation. Philosophy of Science, 1(3):361– 365, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184598.

M:1934:CPC

[41] W. M. M. Classification of paradoxers (continued). Philosophy of Science, 1(3):366–367, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184599.

Anonymous:1934:BMc

[42] Anonymous. Back matter. Philosophy of Science, 1(3):??, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184600.

Anonymous:1934:FMc

[43] Anonymous. Front matter. Philosophy of Science, 1(3):??, July 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184588.

Smith:1934:ALS

[44] Henry Bradford Smith. Abstract logic or the science of modality. Philosophy of Science, 1(4):369–397, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184639.

Malisoff:1934:EQTc

[45] William Marias Malisoff. An examination of the quantum theories III. Philosophy of Science, 1(4):398–408, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184640.

Rashevsky:1934:PMA

[46] N. Rashevsky. Physico-mathematical aspects of the gestalt-problem. Philosophy of Science, 1(4):409–419, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184641.

Feigl:1934:LAP

[47] Herbert Feigl. Logical analysis of the psychophysical problem: A contribution of the new positivism. Philosophy of Science, 1(4):420–445, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184642.

Hartshorne:1934:PDM

[48] Charles Hartshorne. The parallel development of method in physics and psychology. Philosophy of Science, 1(4):446–459, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184643.

Norris:1934:POT

[49] O. O. Norris. A preamble to an organismic theory of knowledge. Philosophy of Science, 1(4):460–478, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184644.

Wiener:1934:QMH

[50] N. Wiener. Quantum mechanics, Haldane, and Leibniz. Philosophy of Science, 1(4):479–482, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184645.

Dubs:1934:PI

[51] Homer H. Dubs and Herbert Feigl. The principle of induction. Philosophy of Science, 1(4):482–486, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184646.

Margenau:1934:FST

[52] Henry Margenau. Flexibility of scientific truth. Philosophy of Science, 1 (4):486–487, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184647.

Smith:1934:BRBb

[53] Henry Bradford Smith. Book review: An Introduction to Logic and Scientific Method by Morris R. Cohen; Ernest Nagel. Philosophy of Science, 1(4):488–491, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184648.

M:1934:NCB

[54] W. M. M. Notes from a case-book. Philosophy of Science, 1(4):492–493, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184649.

Anonymous:1934:BMd

[55] Anonymous. Back matter. Philosophy of Science, 1(4):??, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184650.

Anonymous:1934:FMd

[56] Anonymous. Front matter. Philosophy of Science, 1(4):??, October 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184638.

Blumberg:1935:NPA

[57] Albert E. Blumberg. The nature of philosophic analysis. Philosophy of Science, 2(1):1–8, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184553.

Emery:1935:DVM

[58] A. Emery. Dialectics versus mechanics. A communist debate on scientific method. Philosophy of Science, 2(1):9–38, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184554.

Franklin:1935:WT

[59] Philip Franklin. What is topology? Philosophy of Science, 2(1):39–47, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184555.

Margenau:1935:MMPa

[60] Henry Margenau. Methodology of modern physics. Philosophy of Science, 2(1):48–72, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184556.

Rashevsky:1935:BST

[61] N. Rashevsky. The biophysics of space and time. Philosophy of Science, 2(1):73–85, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184557.

Bliss:1935:SSO

[62] Henry Evelyn Bliss. The system of the sciences and the organization of knowledge. Philosophy of Science, 2(1):86–103, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184558.

Fraenkel:1935:CMN

[63] Adolf Fraenkel and E. T. Bell. Concerning the method of numberpairs. Philosophy of Science, 2(1):104–111, January 1935. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184559.

Henle:1935:NVA

[64] Paul Henle and Henry Bradford Smith. A note on the validity of Aristotelian logic. Philosophy of Science, 2(1):111–114, January 1935. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184560. See [17].

Bell:1935:BRT

[65] E. T. Bell. Book review: The troubled waters of mathematics: The Nature of Mathematics. A critical survey by Max Black. Philosophy of Science, 2(1):115–117, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184561.

Reiser:1935:BRB

[66] Oliver L. Reiser. Book review: The Search for Truth by Eric T. Bell. Philosophy of Science, 2(1):118–120, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184562.

Anonymous:1935:FMa

[67] Anonymous. Front matter. Philosophy of Science, 2(1):??, January 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184552.

Ducasse:1935:SVP

[68] C. J. Ducasse. Is scientific verification possible in philosophy? Philosophy of Science, 2(2):121–127, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184779.

Schilpp:1935:NG

[69] Paul Arthur Schilpp. The nature of the ‘given’. Philosophy of Science, 2(2):128–138, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184780.

Kattsoff:1935:PM

[70] Louis Osgood Kattsoff. Postulational methods. I. Philosophy of Science, 2(2):139–163, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184781.

Margenau:1935:MMPb

[71] Henry Margenau. Methodology of modern physics. Philosophy of Science, 2(2):164–187, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184782.

Ruja:1935:SAC

[72] Harry Ruja. Samuel Alexander’s concept of space-time. Philosophy of Science, 2(2):188–209, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184783.

Frank:1935:SFG

[73] Lawrence K. Frank. Structure, function and growth. Philosophy of Science, 2(2):210–235, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184784.

Boring:1935:RAS

[74] Edwin G. Boring. The relation of the attributes of sensation to the dimensions of the stimulus. Philosophy of Science, 2(2):236–245, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184785.

Somerville:1935:FPA

[75] John Somerville. The forgotten problem: Aims in science. Philosophy of Science, 2(2):246–254, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184786.

Hartshorne:1935:FST

[76] Charles Hartshorne. Flexibility of scientific truth. Philosophy of Science, 2(2):255–256, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184787.

Aldrich:1935:SLP

[77] V. C. Aldrich and Herbert Feigl. Spatial location and the psycho-physical problem. Philosophy of Science, 2(2):256–261, April 1935. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184788.

Witmer:1935:BRB

[78] Enos E. Witmer. Book review: Relativity, Thermodynamics and Cosmology by Richard C. Tolman. Philosophy of Science, 2(2):262–265, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184789.

Clarke:1935:BRB

[79] Francis P. Clarke. Book review: Aristotle’s Theory of the Infinite by Abraham Edel. Philosophy of Science, 2(2):266, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184790.

Malisoff:1935:BRBa

[80] Harry Malisoff. Book review: The Two Poles of the Mind by G. Mannoury. Philosophy of Science, 2(2):267–268, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184791.

Anonymous:1935:BRR

[81] Anonymous. Books received for review. Philosophy of Science, 2(2):268, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184792.

Anonymous:1935:Aa

[82] Anonymous. Announcement. Philosophy of Science, 2(2):268–269, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184793.

Anonymous:1935:NCB

[83] Anonymous. Notes from a case-book. Philosophy of Science, 2(2):270, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184794.

Anonymous:1935:FMb

[84] Anonymous. Front matter. Philosophy of Science, 2(2):??, April 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184778.

Morris:1935:PSS

[85] Charles W. Morris. Philosophy of science and science of philosophy. Philosophy of Science, 2(3):271–286, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184524.

Hartshorne:1935:MP

[86] Charles Hartshorne. Metaphysics for positivists. Philosophy of Science, 2(3):287–303, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184525.

Birkhoff:1935:SCS

[87] G. D. Birkhoff and D. C. Lewis, Jr. Stability in causal systems. Philosophy of Science, 2(3):304–333, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184526.

Malisoff:1935:EQT

[88] William Marias Malisoff. An examination of the quantum theories. IV. Philosophy of Science, 2(3):334–343, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184527.

Dobzhansky:1935:CSC

[89] Th. Dobzhansky. A critique of the species concept in biology. Philosophy of Science, 2(3):344–355, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184528.

Tolman:1935:PVI

[90] Edward Chace Tolman. Psychology versus immediate experience. Philosophy of Science, 2(3):356–380, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184529.

Malisoff:1935:BRBb

[91] W. M. Malisoff. Book review: Philosophy of Science by Fulton J. Sheen. Philosophy of Science, 2(3):381–383, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184530.

M:1935:BRB

[92] W. M. M. Book review: Science and the Public Mind by Benjamin C. Gruenberg. Philosophy of Science, 2(3):383, July 1935. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184531.

Smith:1935:BRB

[93] Henry Bradford Smith. Book review: A System of Logistic by Willard van Orman Quine. Philosophy of Science, 2(3):383–384, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184532.

Anonymous:1935:BR

[94] Anonymous. Books received. Philosophy of Science, 2(3):384, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184533.

Anonymous:1935:Ab

[95] Anonymous. Announcements. Philosophy of Science, 2(3):384–385, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184534.

M:1935:NCB

[96] W. M. M. Notes from a case-book. Philosophy of Science, 2(3):386, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184535.

Anonymous:1935:FMc

[97] Anonymous. Front matter. Philosophy of Science, 2(3):??, July 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184523.

Weinberg:1935:TUS

[98] Julius R. Weinberg. Are there ultimate simples? Philosophy of Science, 2(4):387–399, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184437.

Kolman:1935:PUP

[99] E. Kolman. On the problem of a unified physical theory of matter. Philosophy of Science, 2(4):400–412, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184438.

Rashevsky:1935:OMT

[100] N. Rashevsky. Outline of a mathematical theory of human relations. Philosophy of Science, 2(4):413–430, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184439.

Mayer:1935:TBC

[101] Joseph Mayer. The techniques, basic concepts, and preconceptions of science and their relation to social study. Philosophy of Science, 2(4):431– 483, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184440.

Cairns:1935:LSS

[102] Huntington Cairns. Law as a social science. Philosophy of Science, 2 (4):484–498, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184441.

Bloomfield:1935:LAS

[103] Leonard Bloomfield. Linguistic aspects of science. Philosophy of Science, 2(4):499–517, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184442.

Muenzinger:1935:MVO

[104] Karl F. Muenzinger. Mechanism, vitalism and the organismic hypothesis. Philosophy of Science, 2(4):518–520, October 1935. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184443.

Anonymous:1935:VI

[105] Anonymous. Volume information. Philosophy of Science, 2(4):521–523, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184435.

Anonymous:1935:FMd

[106] Anonymous. Front matter. Philosophy of Science, 2(4):??, October 1935. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184436.

Moore:1936:MDP

[107] Merritt Hadden Moore. A metaphysics of design without purpose. Philosophy of Science, 3(1):1–8, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184752.

Lapan:1936:STA

[108] Arthur Lapan. On space and time as attributes of nature and forms of experience. Philosophy of Science, 3(1):9–18, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184753.

Miller:1936:TP

[109] David L. Miller. Things and potentiality. Philosophy of Science, 3(1):19– 25, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184754.

Wheeler:1936:OLH

[110] Raymond Holder Wheeler. Organismic logic in the history of science. Philosophy of Science, 3(1):26–61, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184755.

Nelson:1936:NP

[111] Everett J. Nelson. A note on parsimony. Philosophy of Science, 3(1):62– 66, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184756.

Kattsoff:1936:PMIa

[112] Louis Osgood Kattsoff. Postulational methods. II. Philosophy of Science, 3(1):67–89, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184757.

Stevens:1936:PPS

[113] S. S. Stevens. Psychology: The propaedeutic science. Philosophy of Science, 3(1):90–103, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184758.

Ratner:1936:ERS

[114] Sidney Ratner. Evolution and the rise of the scientific spirit in America. Philosophy of Science, 3(1):104–122, January 1936. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184759.

Smith:1936:LT

[115] H. B. Smith. The law of transitivity. Philosophy of Science, 3(1):123, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184760.

Reichenbach:1936:IP

[116] Hans Reichenbach. Induction and probability. Philosophy of Science, 3 (1):124–126, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184761.

Morris:1936:BRBa

[117] Charles W. Morris. Book review: Einheit der Wissenschaft. Prager Vorkonferenz der Internationalen Kongresse Fuer Einheit der Wissenschaft — 1934. Philosophy of Science, 3(1):127–130, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184762.

Malisoff:1936:BRBa

[118] Harry Malisoff. Book review: La Logique et sa Caricature by Marcel Boll. Philosophy of Science, 3(1):130–131, January 1936. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184763.

Anonymous:1936:BRa

[119] Anonymous. Books received. Philosophy of Science, 3(1):131, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184764.

Anonymous:1936:FMa

[120] Anonymous. Front matter. Philosophy of Science, 3(1):??, January 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184751.

Gaviola:1936:IIB

[121] E. Gaviola. The impossibility of interaction between mind and matter. Philosophy of Science, 3(2):133–142, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184340.

Fries:1936:ECM

[122] Horace S. Fries. On an empirical criterion of meaning. Philosophy of Science, 3(2):143–151, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184341.

Abraham:1936:NFD

[123] Leo Abraham. A note on the fruitfulness of deduction. Philosophy of Science, 3(2):152–155, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184342.

Baylis:1936:SPN

[124] Charles A. Baylis. Are some propositions neither true nor false? Philosophy of Science, 3(2):156–166, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184343.

Wohlstetter:1936:SPF

[125] Albert Wohlstetter. The structure of the proposition and the fact. Philosophy of Science, 3(2):167–184, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184344.

Emch:1936:CIP

[126] Arnold F. Emch. Consistency and Independence in postulational technique. Philosophy of Science, 3(2):185–196, April 1936. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184345.

Bell:1936:DKP

[127] E. T. Bell. A detail in Kronecker’s program. Philosophy of Science, 3(2): 197–207, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184346.

Irving:1936:LTM

[128] J. A. Irving. Leibniz’ theory of matter. Philosophy of Science, 3(2):208– 214, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184347.

Northrop:1936:PSC

[129] F. S. C. Northrop. The philosophical significance of the concept of probability in quantum mechanics. Philosophy of Science, 3(2):215–232, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184348.

Polanyi:1936:VI

[130] M. Polanyi. The value of the inexact. Philosophy of Science, 3(2):233– 234, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184349.

Jones:1936:IA

[131] Philip C. Jones. Immunity of the apriori. Philosophy of Science, 3(2):234– 235, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184350.

Anonymous:1936:BRb

[132] Anonymous. Books received. Philosophy of Science, 3(2):236, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184351.

Anonymous:1936:FMb

[133] Anonymous. Front matter. Philosophy of Science, 3(2):??, April 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184339.

Nelson:1936:IAE

[134] Everett J. Nelson. The inductive argument for an external world. Philosophy of Science, 3(3):237–249, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184664.

Benjamin:1936:OET

[135] A. Cornelius Benjamin. Outlines of an empirical theory of meaning. Philosophy of Science, 3(3):250–266, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/184665. See comment [164].

Miller:1936:PDP

[136] David L. Miller. Purpose, design and physical relativity. Philosophy of Science, 3(3):267–285, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184666.

Norris:1936:LST

[137] O. O. Norris. The logic of science and technology. Philosophy of Science, 3(3):286–306, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184667.

Wiener:1936:RO

[138] Norbert Wiener. The role of the observer. Philosophy of Science, 3(3):307– 319, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184668.

Gruen:1936:QLU

[139] William Gruen. The quest for a lost universe. Philosophy of Science, 3 (3):320–333, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184669.

Mayer:1936:PSEa

[140] Joseph Mayer. Pseudo-scientific economic doctrine. Philosophy of Science, 3(3):334–359, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184670.

Malisoff:1936:BRU

[141] William Marias Malisoff. Book review: The universe of operations: The Nature of Physical Theory, by P. W. Bridgman. Philosophy of Science, 3 (3):360–364, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184671.

Watson:1936:BRH

[142] David Lindsay Watson. Book review: Humanizing science: Philosophy and the Concepts of Modern Science, by Oliver L. Reiser. Philosophy of Science, 3(3):364–365, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184672.

Lowe:1936:BRT

[143] Victor Lowe. Book review: A tribute to Whitehead: Philosophical Essays for Alfred North Whitehead. Philosophy of Science, 3(3):366–370, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184673.

Malisoff:1936:BRS

[144] William Marias Malisoff. Book review: Solid fare of physics: Foundations of Physics. Philosophy of Science, 3(3):371–372, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184674.

Anonymous:1936:BRc

[145] Anonymous. Books received. Philosophy of Science, 3(3):372–373, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184675.

Kattsoff:1936:PMIb

[146] Louis Osgood Kattsoff. Postulational methods. III. Philosophy of Science, 3(3):375–417, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184676.

Anonymous:1936:FMc

[147] Anonymous. Front matter. Philosophy of Science, 3(3):??, July 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184663.

Carnap:1936:TM

[148] Rudolf Carnap. Testability and meaning. Philosophy of Science, 3(4):419– 471, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184400.

Bentley:1936:PL

[149] Arthur F. Bentley. The positive and the logical. Philosophy of Science, 3 (4):472–485, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184401.

Fitch:1936:PC

[150] Frederic B. Fitch. Physical continuity. Philosophy of Science, 3(4):486– 493, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184402.

Birch:1936:TCW

[151] T. Bruce Birch. The theory of continuity of William of Ockham. Philosophy of Science, 3(4):494–505, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184403.

Aldrich:1936:RI

[152] V. C. Aldrich. Renegade instances. Philosophy of Science, 3(4):506–514, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184404.

Mayer:1936:PSEb

[153] Joseph Mayer. Pseudo-scientific economic doctrine — continued. Philosophy of Science, 3(4):515–541, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184405.

Morris:1936:BRBb

[154] Charles W. Morris. Book review: Le Developpement du Cercle de Vienne et l’Avenir de l’Empirisme Logique by Otto Neurath. Philosophy of Science, 3(4):542–543, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184406.

Reiser:1936:BRC

[155] Oliver L. Reiser. Book review: Chance is real: A World of Chance. Philosophy of Science, 3(4):543–545, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184407.

Reiser:1936:BRA

[156] Oliver L. Reiser. Book review: Aristotle, Galileo and the leaning towers of science: Principles of Topological Psychology by Kurt Lewin. Philosophy of Science, 3(4):545–548, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184408.

Anonymous:1936:BRd

[157] Anonymous. Books received. Philosophy of Science, 3(4):548–549, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184409.

Smith:1936:AP

[158] H. B. Smith. The algebra of propositions. Philosophy of Science, 3(4): 551–578, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184410.

Anonymous:1936:FMd

[159] Anonymous. Front matter. Philosophy of Science, 3(4):??, October 1936. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184399.

Carnap:1937:TMC

[160] Rudolf Carnap. Testability and meaning — continued. Philosophy of Science, 4(1):1–40, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184580.

Frank:1937:MVM

[161] Philipp Frank and Philip Shorr. The mechanical versus the mathematical conception of nature. Philosophy of Science, 4(1):41–74, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184581.

Chalmers:1937:LUM

[162] Gordon Keith Chalmers. The lodestone and the understanding of matter in Seventeenth Century England. Philosophy of Science, 4(1):75–95, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184582.

Rosenzweig:1937:SPC

[163] Saul Rosenzweig. Schools of psychology: A complementary pattern. Philosophy of Science, 4(1):96–106, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184583.

Biederman:1937:COE

[164] Israel C. Biederman. [comment on “outlines of an empirical theory of meaning”]. Philosophy of Science, 4(1):107–108, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184584. See [135].

Leonard:1937:BRP

[165] Henry S. Leonard. Book review: The pragmatism and scientific metaphysics of C. S. Peirce: Collected Papers of C. S. Peirce. Vol. V. Pragmatism and Pragmaticism by Charles Sanders Peirce; Charles Hartshorne; Paul Weiss; Collected Papers of C. S. Peirce. Vol. VI. Scientific Metaphysics by Charles Sanders Peirce; Charles Hartshorne; Paul Weiss. Philosophy of Science, 4(1):109–121, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184585.

M:1937:NCB

[166] W. M. M. Notes from a case-book. Philosophy of Science, 4(1):122–123, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184586.

Hailperin:1937:FPM

[167] Theodore Hailperin. Foundations of probability in mathematical logic. Philosophy of Science, 4(1):125–150, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/184587.

Anonymous:1937:FMa

[168] Anonymous. Front matter. Philosophy of Science, 4(1):??, January 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184579.

Lindsay:1937:MSP

[169] R. B. Lindsay. The meaning of simplicity in physics. Philosophy of Science, 4(2):151–167, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184860.

Senior:1937:SDC

[170] James K. Senior. Some dangers connected with mathematical applications. Philosophy of Science, 4(2):168–172, April 1937. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184861.

Richmond:1937:MTF

[171] Carl A. Richmond. The measurement of time: A first chapter of physics. Philosophy of Science, 4(2):173–201, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184862.

Lillie:1937:DAL

[172] Ralph S. Lillie. Directive action and life. Philosophy of Science, 4(2):202– 226, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184863.

Brunswik:1937:PSO

[173] Egon Brunswik. Psychology as a science of objective relations. Philosophy of Science, 4(2):227–260, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184864.

Malisoff:1937:ASE

[174] William Marias Malisoff. Arranging the sciences: I. An experiment. Philosophy of Science, 4(2):261–264, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184865.

Neurath:1937:USE

[175] Otto Neurath. Unified science and its encyclopaedia. Philosophy of Science, 4(2):265–277, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184866.

Smith:1937:LML

[176] Henry Bradford Smith. On logics that are more or less abstract. Philosophy of Science, 4(2):278–281, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184867.

Benjamin:1937:MM

[177] A. Cornelius Benjamin. The meaning of meaning. Philosophy of Science, 4(2):282, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184868.

Roberts:1937:NCS

[178] John L. Roberts. A new classification of statistics. Philosophy of Science, 4(2):283–284, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184869.

Konvitz:1937:BRB

[179] Milton R. Konvitz. Book review: An Examination of Logical Positivism by Julius Rudolph Weinberg. Philosophy of Science, 4(2):285–287, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184870.

Anonymous:1937:BR

[180] Anonymous. Books received. Philosophy of Science, 4(2):287, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184872.

Anonymous:1937:CHF

[181] Anonymous. Corrections to Hailperin’s “Foundations of Probability in Mathematical Logic.” Supplement, January, 1937. Philosophy of Science, 4(2):287, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184873.

M:1937:BRB

[182] W. M. M. Book review: The Einstein Theory of Relativity by H. G. L. R. Lieber. Philosophy of Science, 4(2):287, April 1937. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184871.

Anonymous:1937:FMb

[183] Anonymous. Front matter. Philosophy of Science, 4(2):??, April 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184859.

Bohr:1937:CC

[184] Niels Bohr. Causality and complementarity. Philosophy of Science, 4(3): 289–298, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184445.

Menger:1937:NL

[185] Karl Menger, H. B. Gottlieb, and J. K. Senior. The new logic. Philosophy of Science, 4(3):299–336, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184446.

Margenau:1937:CPM

[186] Henry Margenau. Critical points in modern physical theory. Philosophy of Science, 4(3):337–370, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184447.

Bisbee:1937:OSS

[187] Eleanor Bisbee. Objectivity in the social sciences. Philosophy of Science, 4(3):371–382, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184448.

Smith:1937:MLR

[188] H. B. Smith. Modal logic — a revision. Philosophy of Science, 4(3):383– 384, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184449.

Malisoff:1937:BRA

[189] William Marias Malisoff. Book review: The anatomy of science: The Logical Structure of Science. Philosophy of Science, 4(3):385–387, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184450.

Wiener:1937:BRP

[190] Philip Paul Wiener. Book review: Philosophical history: Histoire de la Pensee Scientifique. Philosophy of Science, 4(3):387–390, July 1937. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184451.

Hofstadter:1937:BRE

[191] Albert Hofstadter. Book review: Empirical probability: Wahrscheinlichkeit Statistik und Wahrheit, by Richard von Mises. Einfuhrung in Die Neue Wahrscheinlichkeitslehre und Ihre Anwendung. Zweite, Neubearbeitete Auflage. Philosophy of Science, 4(3):390–392, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184452.

Reiser:1937:BRM

[192] Oliver L. Reiser. Book review: A memory continuum? The Psychology of Belief. Philosophy of Science, 4(3):392–393, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184453.

Anonymous:1937:FMc

[193] Anonymous. Front matter. Philosophy of Science, 4(3):??, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184444.

Aldrich:1937:DMD

[194] Virgil C. Aldrich. DesCartes’ method of doubt. Philosophy of Science, 4 (4):395–411, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184412.

Tugwell:1937:TX

[195] R. G. Tugwell. On the troublesome “x”. Philosophy of Science, 4(4):412– 426, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184413.

Black:1937:VEL

[196] Max Black. Vagueness. An exercise in logical analysis. Philosophy of Science, 4(4):427–455, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184414.

Lindsay:1937:COP

[197] R. B. Lindsay. A critique of operationalism in physics. Philosophy of Science, 4(4):456–470, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184415.

Sisson:1937:SLE

[198] Edward O. Sisson. Symbolic logic and “embedding language”. Philosophy of Science, 4(4):471–481, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184416.

Huntington:1937:MP

[199] Edward V. Huntington. The method of postulates. Philosophy of Science, 4(4):482–495, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184417.

Anonymous:1937:SUS

[200] Anonymous. Symposium of unified science. Philosophy of Science, 4(4): 496–498, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184418.

Reiser:1937:BRY

[201] Oliver L. Reiser. Book review: Yoga: Yoga: A Scientific Evaluation by Kovoor T. Behanan. Philosophy of Science, 4(4):499, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184419.

Anonymous:1937:FMd

[202] Anonymous. Front matter. Philosophy of Science, 4(4):??, October 1937. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184411.

Bures:1938:CPC

[203] Charles E. Bures. The concept of probability. A critical survey of recent contributions. Philosophy of Science, 5(1):1–20, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184701.

Reichenbach:1938:PI

[204] Hans Reichenbach. On probability and induction. Philosophy of Science, 5(1):21–45, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184702.

Nagel:1938:STP

[205] Ernest Nagel. Some theses in the philosophy of logic. Philosophy of Science, 5(1):46–51, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184703.

Skulsky:1938:TT

[206] Samuel Skulsky. A theory of time. Philosophy of Science, 5(1):52–59, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184704.

Dotterer:1938:I

[207] Ray H. Dotterer. Indeterminisms. Philosophy of Science, 5(1):60–72, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184705.

Hillman:1938:EMS

[208] Owen N. Hillman. Emile Meyerson on scientific explanation. Philosophy of Science, 5(1):73–80, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184706.

Moore:1938:SSF

[209] Willis Moore. Structure in sentence and in fact. Philosophy of Science, 5(1):81–88, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184707.

Lee:1938:CII

[210] D. Demetracopoulou Lee. Conceptual implications of an Indian language. Philosophy of Science, 5(1):89–102, January 1938. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184708.

Dotterer:1938:CIP

[211] Ray H. Dotterer. The conception of infinite progress. Philosophy of Science, 5(1):103–104, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184709.

Brunswik:1938:BRB

[212] Egon Brunswik. Book review: Erkenntnis und Wissenschaftliches Verhalten by Arne Ness. Philosophy of Science, 5(1):105–107, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184710.

Reiser:1938:BRT

[213] Oliver L. Reiser. Book review: Time and tide will wait: The Span of Life. Philosophy of Science, 5(1):107–108, January 1938. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184711.

Anonymous:1938:BRa

[214] Anonymous. Books received. Philosophy of Science, 5(1):108–109, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184712.

Anonymous:1938:EPS

[215] Anonymous. Errata: Psychology as a science of objective relations. Philosophy of Science, 5(1):110, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184713.

Anonymous:1938:FMa

[216] Anonymous. Front matter. Philosophy of Science, 5(1):??, January 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184700.

M:1938:EWP

[217] W. M. M. Editorial: Where is philosophy of science? Philosophy of Science, 5(2):111–113, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184625.

Bridgman:1938:OA

[218] P. W. Bridgman. Operational analysis. Philosophy of Science, 5(2):114– 131, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184626.

Bentley:1938:PF

[219] Arthur F. Bentley. Physicists and fairies. Philosophy of Science, 5(2):132– 165, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184627.

Northrop:1938:CFP

[220] F. S. C. Northrop. Causality in field physics in its bearing upon biological causation. Philosophy of Science, 5(2):166–180, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184628.

Williams:1938:IEW

[221] Donald C. Williams. Induction and the external world. Philosophy of Science, 5(2):181–188, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184629.

Lundberg:1938:CLS

[222] George A. Lundberg. The concept of law in the social sciences. Philosophy of Science, 5(2):189–203, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184630.

Bierstedt:1938:MC

[223] Robert Bierstedt. The meanings of culture. Philosophy of Science, 5(2): 204–216, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184631.

Lowe:1938:MMI

[224] Victor Lowe. Mr. Miller’s interpretation of Whitehead. Philosophy of Science, 5(2):217–229, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184632.

Nagel:1938:BRBa

[225] Ernest Nagel. Book review: Determinismus und Indeterminismus in der Modernen Physik by Ernst Cassirer. Philosophy of Science, 5(2):230– 232, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184633.

Nagel:1938:BRBb

[226] Ernest Nagel. Book review: Einfuhrung in Die Mathematische Logik by Alfred Tarski. Philosophy of Science, 5(2):232–233, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184634.

Malisoff:1938:BRB

[227] Harry Malisoff. Book reviews: Time’s Arrow in Society by Anderson Woods; Recent Theories of Sovereignty by Hyman Ezra Cohen. Philosophy of Science, 5(2):233–234, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184635.

Anonymous:1938:FIC

[228] Anonymous. Fourth international Congress for the unity of science. Philosophy of Science, 5(2):234–235, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184636.

Anonymous:1938:BRb

[229] Anonymous. Books received. Philosophy of Science, 5(2):235, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184637.

Anonymous:1938:FMb

[230] Anonymous. Front matter. Philosophy of Science, 5(2):??, April 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184624.

Prescott:1938:SME

[231] C. H. Prescott, Jr. The scientific method and its extension to systems of many degrees of freedom. Philosophy of Science, 5(3):237–266, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184833.

Boodin:1938:RMS

[232] John Elof Boodin. A revolution in metaphysics and in science. Philosophy of Science, 5(3):267–275, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184834.

Urban:1938:SSP

[233] Wilbur M. Urban. Symbolism in science and philosophy. Philosophy of Science, 5(3):276–299, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184835.

Vivas:1938:NCS

[234] Eliseo Vivas. Nature common sense and science. Philosophy of Science, 5(3):300–309, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184836.

Lepley:1938:VFV

[235] Ray Lepley. The verifiability of facts and values. Philosophy of Science, 5(3):310–320, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184837.

Merton:1938:SSO

[236] Robert K. Merton. Science and the social order. Philosophy of Science, 5(3):321–337, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184838.

Harding:1938:STB

[237] T. Swann Harding. Science at the tower of Babel. Philosophy of Science, 5(3):338–353, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184839.

Metcalf:1938:IEW

[238] Wilmot V. Metcalf. Induction of an external world. Philosophy of Science, 5(3):354–358, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184840.

Samachson:1938:BRO

[239] Joseph Samachson. Book review: The old god of physics: Isaac Newton. Philosophy of Science, 5(3):359–360, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184841.

Bisbee:1938:BRW

[240] Eleanor Bisbee. Book review: A world of probability: Experience and Prediction. Philosophy of Science, 5(3):360–366, July 1938. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184842.

Anonymous:1938:BRc

[241] Anonymous. Books received. Philosophy of Science, 5(3):367–368, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184843.

Anonymous:1938:FMc

[242] Anonymous. Front matter. Philosophy of Science, 5(3):??, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184832.

White:1938:SS

[243] Leslie A. White. Science is sciencing. Philosophy of Science, 5(4):369–389, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184653.

Malisoff:1938:ASI

[244] William Marias Malisoff. Arranging the sciences. II. another experiment. Philosophy of Science, 5(4):390–392, October 1938. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184654.

Lapan:1938:PTN

[245] Arthur Lapan. Preface to a theory of nature. Philosophy of Science, 5 (4):393–409, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184655.

Taube:1938:RES

[246] Mortimer Taube. A re-examination of some arguments for realism. Philosophy of Science, 5(4):410–420, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184656.

Benjamin:1938:SPS

[247] A. C. Benjamin. Science and the philosophy of science. Philosophy of Science, 5(4):421–433, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184657.

Jeffreys:1938:NM

[248] Harold Jeffreys. The nature of mathematics. Philosophy of Science, 5 (4):434–451, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184658.

Lindsay:1938:FTP

[249] R. B. Lindsay. The future of theoretical physics. Philosophy of Science, 5(4):452–471, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184659.

Rogosin:1938:TPR

[250] H. Rogosin. Telepathy, psychical research, and modern psychology. Philosophy of Science, 5(4):472–483, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184660.

Neurath:1938:EPA

[251] Otto Neurath. Encyclopaedism as a pedagogical aim: A Danish approach. Philosophy of Science, 5(4):484–492, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184661.

Whitrow:1938:RH

[252] G. J. Whitrow. Robert Hooke. Philosophy of Science, 5(4):493–502, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184662.

Anonymous:1938:VI

[253] Anonymous. Volume information. Philosophy of Science, 5(4):503–505, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184651.

Anonymous:1938:FMd

[254] Anonymous. Front matter. Philosophy of Science, 5(4):??, October 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184652.

Ablowitz:1939:TE

[255] Reuben Ablowitz. The theory of emergence. Philosophy of Science, 6(1): 1–16, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184327.

Malisoff:1939:EM

[256] William Marias Malisoff. Emergence without mystery. Philosophy of Science, 6(1):17–24, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184328.

Angyal:1939:SW

[257] Andras Angyal. The structure of wholes. Philosophy of Science, 6(1):25– 37, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184329.

Peddie:1939:PIP

[258] W. Peddie. The philosophy of “as if” in physical science. Philosophy of Science, 6(1):38–47, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184330.

Darwin:1939:LPP

[259] C. G. Darwin. Logic and probability in physics. Philosophy of Science, 6(1):48–64, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184331.

Margenau:1939:PMV

[260] Henry Margenau. Probability, many-valued logics, and physics. Philosophy of Science, 6(1):65–87, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184332.

Ratner:1939:PCH

[261] Sidney Ratner. Patterns of culture in history. Philosophy of Science, 6 (1):88–97, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184333.

Harding:1939:MPR

[262] T. Swann Harding. The mass production of research. Philosophy of Science, 6(1):98–105, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184334.

Bross:1939:RCC

[263] John R. Bross and George J. Bowdery. A realistic criticism of a contemporary philosophy of logic. Philosophy of Science, 6(1):105–114, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184335.

Nathanson:1939:BRD

[264] Jerome Nathanson. Book review: Dewey’s vivisection of the logical process: Logic: The Theory of Inquiry. Philosophy of Science, 6(1):115–122, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184336.

Anonymous:1939:BRa

[265] Anonymous. Books received. Philosophy of Science, 6(1):122–125, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184337.

Anonymous:1939:Aa

[266] Anonymous. Announcement. Philosophy of Science, 6(1):126, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184338.

Anonymous:1939:FMa

[267] Anonymous. Front matter. Philosophy of Science, 6(1):??, January 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184326.

Malisoff:1939:VS

[268] William Marias Malisoff. Virtue and the scientist. Philosophy of Science, 6(2):127–136, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184564.

McCrea:1939:ETS

[269] W. H. McCrea. The evolution of theories of space-time and mechanics. Philosophy of Science, 6(2):137–162, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184565.

Hempel:1939:VL

[270] Carl G. Hempel. Vagueness and logic. Philosophy of Science, 6(2):163– 180, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184566.

Copilowish:1939:BLC

[271] Irving M. Copilowish. Border-line cases, vagueness, and ambiguity. Philosophy of Science, 6(2):181–195, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184567.

McGill:1939:CLC

[272] V. J. McGill. Concerning the laws of contradiction and excluded middle. Philosophy of Science, 6(2):196–211, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184568.

Nagel:1939:PTK

[273] Ernest Nagel. Probability and the theory of knowledge. Philosophy of Science, 6(2):212–253, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184569.

Lenzen:1939:HD

[274] V. F. Lenzen. The hypothesis of dualism. Philosophy of Science, 6(2):254– 256, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184570.

Harms:1939:BRBa

[275] Ernst Harms. Book review: Denker der Italienischen Renaissance. Gestalten und Probleme. (Thinkers of the Italian Renaissance. Figures and Problems.) by Richard Hoenigswald. Philosophy of Science, 6(2):257, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184573.

M:1939:BRBa

[276] W. M. M. Book review: Research and Statistical Methodology Books and Reviews of 1933–1938 by O. K. Buros. Philosophy of Science, 6(2):257, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184571.

M:1939:BRBb

[277] W. M. M. Book review: An Introduction to the History of History by James J. Shotwell. Philosophy of Science, 6(2):257, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184572.

Biser:1939:BRBa

[278] Irwin Biser. Book review: Drei Vortrage Uber Philosophie und Parapschologie by Johannes J. Poortman. Philosophy of Science, 6(2):258, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184575.

Harms:1939:BRBb

[279] Ernst Harms. Book review: Geist und Sein. (Mind and Being.) by Herman Schmalenbach. Philosophy of Science, 6(2):258, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184574.

Meisels:1939:BRBa

[280] Karol Meisels. Book review: Critique de la Mesure by G. Beneze. Philosophy of Science, 6(2):258–259, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184576.

Meisels:1939:BRBb

[281] Karol Meisels. Book review: Essai sur les Definitons Experimentales des Operations Chimiques by Paul Renaud. Philosophy of Science, 6(2):259, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184577.

Anonymous:1939:Ab

[282] Anonymous. Announcement. Philosophy of Science, 6(2):260, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184578.

Anonymous:1939:FMb

[283] Anonymous. Front matter. Philosophy of Science, 6(2):??, April 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184563.

Malisoff:1939:WA

[284] William Marias Malisoff. What is an atom? Philosophy of Science, 6(3): 261–265, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184721.

Seifriz:1939:MIL

[285] William Seifriz. A materialistic interpretation of life. Philosophy of Science, 6(3):266–284, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184722.

Winn:1939:NR

[286] Ralph B. Winn. Is nature rational? Philosophy of Science, 6(3):285– 300, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184723.

Fries:1939:PVA

[287] Horace S. Fries. Physics, a vicious abstraction. Philosophy of Science, 6 (3):301–308, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184724.

Hardie:1939:OKO

[288] C. D. Hardie. Our knowledge of other minds. Philosophy of Science, 6 (3):309–317, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184725.

Cook:1939:SNS

[289] Thomas I. Cook. Science: Natural and social. Philosophy of Science, 6 (3):318–327, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184726.

Kluckhohn:1939:PTA

[290] Clyde Kluckhohn. The place of theory in anthropological studies. Philosophy of Science, 6(3):328–344, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184727.

Lesser:1939:RPL

[291] Alexander Lesser. Research procedure and laws of culture. Philosophy of Science, 6(3):345–355, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184728.

Werkmeister:1939:NLC

[292] W. H. Werkmeister. Natural languages as cultural indices. Philosophy of Science, 6(3):356–366, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184729.

Metcalf:1939:RA

[293] Wilmot V. Metcalf. The reality of the atom. Philosophy of Science, 6(3): 367–371, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184730.

Dubs:1939:EIP

[294] Homer H. Dubs. Empirical induction by probability and the external world. Philosophy of Science, 6(3):371, July 1939. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184731.

Williams:1939:PDP

[295] Donald C. Williams. Professor Dubs on the Principle of Indifference. Philosophy of Science, 6(3):371–373, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184732.

Jones:1939:ESP

[296] Philip Chapin Jones. Extra-sensory perception and idealism. Philosophy of Science, 6(3):373–374, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184733.

Kaplan:1939:CMP

[297] Oscar Kaplan. Concerning the mass production of research. Philosophy of Science, 6(3):374–377, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184734.

Malisoff:1939:BRB

[298] William Marias Malisoff. Book review: A General Schema for Natural Systems (Nature Considered as a Function of Types of Selectivity and of Modes of Selection) by Irwin Biser. Philosophy of Science, 6(3):378, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184735.

M:1939:BRBc

[299] W. M. M. Book review: A Preview of the Absolute-Relative Theory by Franklin J. Matchette. Philosophy of Science, 6(3):379, July 1939. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184737.

Reiser:1939:BRB

[300] Oliver L. Reiser. Book review: Toward a Dimensional Realism by Charles M. Perry. Philosophy of Science, 6(3):379, July 1939. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184736.

M:1939:BRBd

[301] W. M. M. Book review: “Truth” as Conceived by Those Who Are Not Professional Philosophers by Arne Ness. Philosophy of Science, 6(3):379– 380, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184738.

Biser:1939:BRBb

[302] Irwin Biser. Book review: Die Enstehung des Lebens Durch Stetige Schopfung by Ignaz Lichtig. Philosophy of Science, 6(3):380, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184739.

M:1939:BRBe

[303] W. M. M. Book review: The Old Faiths Perish by Jay William Hudson. Philosophy of Science, 6(3):380, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184740.

M:1939:BRBf

[304] W. M. M. Book review: Who Was Socrates? by Alban D. Winspear; Tom Silverberg. Philosophy of Science, 6(3):380–381, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184741.

Infeld:1939:BRP

[305] Leopold Infeld. Book review: The Principles of Statistical Mechanics by Richard C. Tolman. Philosophy of Science, 6(3):381, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184744.

M:1939:BRBg

[306] W. M. M. Book review: Perpetual Peace by Immanuel Kant. Philosophy of Science, 6(3):381, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184742.

M:1939:BRBh

[307] W. M. M. Book review: Formal Logic by Albert A. Bennett; Charles A. Baylis. Philosophy of Science, 6(3):381, July 1939. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184743.

M:1939:BRBi

[308] W. M. M. Book review: The Concepts of the Calculus by Carl B. Boyer. Philosophy of Science, 6(3):382, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184745.

M:1939:BRBj

[309] W. M. M. Book review: The Scientist in Action by William H. George. Philosophy of Science, 6(3):382, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184746.

M:1939:BRBk

[310] W. M. M. Book review: Crystalline Enzymes by John H. Northrop. Philosophy of Science, 6(3):382, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184747.

M:1939:BRBl

[311] W. M. M. Book review: Chemistry in Relation to Biology and Medicine with Especial Reference to Insulin and Other Hormones by John Jacob Abel. Philosophy of Science, 6(3):382, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184748.

Anonymous:1939:BRb

[312] Anonymous. Books received. Philosophy of Science, 6(3):382–383, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184749.

Anonymous:1939:Ac

[313] Anonymous. Announcement. Philosophy of Science, 6(3):383, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184750.

Anonymous:1939:FMc

[314] Anonymous. Front matter. Philosophy of Science, 6(3):??, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184720.

Malisoff:1939:WG

[315] William Marias Malisoff. What is a gene? Philosophy of Science, 6 (4):385–389, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184494.

Ramsperger:1939:WSK

[316] A. G. Ramsperger. What is scientific knowledge? Philosophy of Science, 6(4):390–403, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184495.

Dexter:1939:CIP

[317] Lewis A. Dexter. Causal imputation and purposes of investigation. Philosophy of Science, 6(4):404–411, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184496.

Sellars:1939:CCR

[318] Roy Wood Sellars. A clarification of critical realism. Philosophy of Science, 6(4):412–421, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184497.

Benjamin:1939:SV

[319] A. Cornelius Benjamin. Science and vagueness. Philosophy of Science, 6 (4):422–431, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184498.

Vivas:1939:VF

[320] Eliseo Vivas. Value and fact. Philosophy of Science, 6(4):432–445, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184499.

Hofstadter:1939:LI

[321] Albert Hofstadter and J. C. C. McKinsey. On the logic of imperatives. Philosophy of Science, 6(4):446–457, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184500.

Nahm:1939:PIS

[322] M. C. Nahm. The philosophical implications of some theories of emotion. Philosophy of Science, 6(4):458–486, October 1939. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184501.

Grelling:1939:CSW

[323] Kurt Grelling and Paul Oppenheim. Concerning the structure of wholes. Philosophy of Science, 6(4):487–489, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/184502.

M:1939:BRBm

[324] W. M. M. Book review: Theory of Valuation by John Dewey. Philosophy of Science, 6(4):490–491, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184503.

M:1939:BRBn

[325] W. M. M. Book review: Beyond Yonder by O. J. Lee. Philosophy of Science, 6(4):491, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184504.

M:1939:BRBo

[326] W. M. M. Book review: La Logique D’Une Idee-Force. L’Idee D’Utilite Sociale et Revolution Francaise (1789–1792) by Jean Belin. Philosophy of Science, 6(4):491, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184505.

Bowdery:1939:BRB

[327] George J. Bowdery and John R. Bross. Book review: Communication by Karl Britton. Philosophy of Science, 6(4):491–493, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184506.

Anonymous:1939:VI

[328] Anonymous. Volume information. Philosophy of Science, 6(4):495–496, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184492.

Anonymous:1939:FMd

[329] Anonymous. Front matter. Philosophy of Science, 6(4):??, October 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184493.